OpenStudy (loser66):

A cylindrical log 15cm in diameter and 68cm long is glowing red hot in a fireplace If it 's emitting radiation at the rate of 38kW, what is its temperature? the log's emissivity e=1 Please , help

OpenStudy (loser66):



OpenStudy (loser66):

\[\large T=\sqrt[4]{\frac{P}{e\sigma A}}\] where P = 3.8*10^4 W e =1 A = \((\dfrac{0.15}{2})^2 * 0.68 = 0.003825 m^2 = 0.3825 *10^{-2}\) am I right?

OpenStudy (loser66):

@LastDayWork

OpenStudy (loser66):

@agent0smith

OpenStudy (loser66):

oh, I calculate A without time to pi :) , now redo :) A = 1.20 x 10^(-2) m^2

OpenStudy (anonymous):

Are you including the areas of both the ends and of the round portion?

OpenStudy (lastdaywork):

My area came out to be 0.782256571 m^2

OpenStudy (loser66):

my calculator said it is 0.355785

OpenStudy (lastdaywork):

What formula are you using to calculate area ?? (in terms of r and h)

OpenStudy (loser66):

area bottom and top are = 2 (r^2 *pi) where r = 0.15/2 = 0.075 area of the side is = 2rpi *h

OpenStudy (lastdaywork):

Okay..then A=0.355785368 m^2 is correct. I only looked at the diagram and got the impression r = 0.15 m XD

OpenStudy (lastdaywork):

Then every thing else is correct; you only need to plug-in the values :)

OpenStudy (loser66):

thank you @LastDayWork I have another question from the concept An 18m by 4.5 m house built on a concrete slab 27 cm thick What is the heat-loss rate throught the floor if the interior is at 29 C while the ground is at 19 C, thermal conductivity of concrete is 1W/m.K I know \(H= -kA (T_f - T_i) in this case, how to define which T is T_f?

OpenStudy (loser66):

and since it is the heat lost (interior temperature > the floor temperature) so, H will have - sign, right?

OpenStudy (anonymous):

Need thickness x there, too heat loss = - k A (T1-T2)/x

OpenStudy (loser66):

oh yes, sorry, I missed it. I calculate and get -3000W . they ask me to write answer using 2 sig fig. is it 30.00 x 10^2 or -30.00 x10^2